OpenStudy (lgbasallote):

is it possible to have double improper integrals? just yes or no....coz i'd like to see how to do them if yes >:))

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

\[\iint \text d x^2=\int x+c~~\text dx=\frac{x^2}2+cx+d\]

OpenStudy (lgbasallote):

lol that's INDEFINITE integral :P \[\int_a^\infty F(x) dx\] that's the improper integral haha

OpenStudy (anonymous):

is i m right lgba...?

OpenStudy (lgbasallote):

double improper let's say \[\int_0^\infty \int_\infty^0 xy dxdy\]

OpenStudy (lgbasallote):

does that happen?

OpenStudy (unklerhaukus):

ah yes @lgbasallote i got my language mixed up

OpenStudy (lgbasallote):

so does it happen?

OpenStudy (unklerhaukus):

would this count \[\int\limits_{-\infty}^{\infty}\text d x\]

OpenStudy (unklerhaukus):

it is doubley improper

OpenStudy (lgbasallote):

lol :P

OpenStudy (lgbasallote):

the one i posted....does it happen?

OpenStudy (unklerhaukus):

i would be surprised if it did not happen, but you limits are upside down on the inner integral

OpenStudy (lgbasallote):

improper integrals =_=

OpenStudy (lgbasallote):

how could it be upside down @UnkleRhaukus i made the problem :P you cant tell the asker what he wants to ask >:)))

OpenStudy (unklerhaukus):

the lowest number on the bottom limit makes more sense as you are integrating from left to right

OpenStudy (lgbasallote):

so bottomline..it's possible?

OpenStudy (anonymous):

Yeah, sure. You could certainly have one. We use improper integrals often when we have asymptotic curves like this: